OpenStudy (anonymous):

i tried a lot of things but i couldnt figure out..Does this series converge ? Σ(n=2to∞) [1/(n-1)(n+1)^1/2+(n+1)(n-1)^1/2 ]

OpenStudy (anonymous):

i wrote the question again for you to see clearly \[\sum_{2}^{\infty} 1/ [(n-1)\sqrt{n+1}+(n+1)\sqrt{n-1}]\]